OpenStudy (anonymous):

For all numbers c and d, let [] be defined by the following equation c [] d = d^2 +cd - c if 3 [] g = 25, what is the positive value of g?

OpenStudy (anonymous):

g=4

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

i guess because where you see a "c"you put 3 and where you see d you put g

OpenStudy (anonymous):

\[3[]g=g^2+3g-3=25\]

myininaya (myininaya):

(g+7)(g-4)=0 g=-7, g=4 g=4 is the only positive solution

OpenStudy (anonymous):

\[g^2+3g-28=0\] \[(g-4)(g+7)=0\] \[g=4\] or \[g=-7\]

OpenStudy (anonymous):

had to beat me didn't you but doesn't count because you didn't use latex nyah nyah

myininaya (myininaya):

yes it does its readable so it counts

OpenStudy (anonymous):

Lol thank you guys for explaining.

myininaya (myininaya):

satellite did all the explaining

OpenStudy (anonymous):

ok you win. now get some sleep. i mean i am going to get some sleep . really. now.